OpenStudy (anonymous):

write the slope-intercept form of the equations of the line described. 1.through (3,-1), parallel to y=3x+4 2.through(-4,-1), parallel to y=-1/2x+3

OpenStudy (anonymous):

y = 3x - 10

OpenStudy (texaschic101):

parallel lines will have the same slope. y = 3x + 4 -- in y = mx + b form, the slope is in the m position. the slope here is 3 (3,-1) x = 3 and y = -1 y = mx + b -1 = 3(3) + b -1 = 9 + b -1 - 9 = b -10 = b parallel equation is : y = 3x - 10 ==================== y = -1/2x + 3 slope = -1/2 (-4,-1) x = -4 and y = -1 y = mx + b -1 = -1/2(-4) + b -1 = 2 + b -1 - 2 = b -3 = b parallel equation is : y = -1/2x - 3